Tech Staff Meeting Notes May 10, 2012
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated May 2012
Arch's retirement dinner is May 22 at 5:30 at Folks Folly.
Duty Table was updated at this staff meeting.
We reviewed the RT software. We want to use trouble tickets instead of mail to ceri_seis (though mail to ceri_seis can still be used).
RDST and MCAR are also prone to hang. They're now on gcf2ew along with a script that forces a kill and restart when necessary. Need a fanless pc.
ISIS station review.
EDIT, NNAR & CHNM: NNAR and CHNM repeat through EDIT and the received signal is very bad due to trees. NNAR and CHRM are very noisy and fade out quite a bit. EDIT is now located in the Chicksaw National Wildlife Preserve so cutting trees is not an option. There isn't a clear path to a repeater on the Edith forest tower either. Need a new plan (maybe go directly to the node?).
WALK: Chris and Manning were unable to reach WALK up in the KY bend. Mr Lynn indicated that the station was completely underwater but it is still up? (yes still up).
Search for quieter site for CTCT.
Dynamic NMDown list
Data taken from the 4.0 Oklahoma event. Here:///gaia/smeagol/dat2/Working/NM/201204/20120403_073430_MAN/
Data taken from: /gaia/smeagol/dat2/Working/NM/201205/20120506_100240_MAN
HCAR - flooded, need to verify that well stays dry.
ARPT.EHN -- gaps in data
Strong Motion station problems
HENM.HNZ - muted amplitude
Mag. 4.0 gain range problems
These stations had problems with the gain channel during the Magnituded 4 Missouri event on Feb 21, 2012. It isn't clear to me whether they data channels saturated or whether there was excessive chatter in the gain channel. See image.
HTAR coax problem.
There were 10 stations down for this event. I'm guessing that for a
majority of them it was just the battery. Chris fixed 4 of them. HCAR was
flooded and HTAR is a likely coax problem. The remaining 4 are listed
LFRT, LVAR, TNMT, and WADM.
Partial strongmo conversion to Basalt started. Needs to be finished.
Jim is getting a quote to replace the generator blockheater, muffler and exhaust manifold, and vaporizor solenoid.
Mitch needs to modify env_alarm to customize for lnxt generator alarms:
Here are what the new Lenox node status alarm channels will be when the new generator interface module is installed. As was the case before, all outputs are normally high and go low to indicate an alarm condition. High is defined as +2.5 VDC or approximately +1024 counts.
- Ch2 Generator Failure, will go low when generator tries to start and fails (was AC power failure)
- Ch3 XLAK GPS IRIG time code generator alarm (no change)
- Ch4 Generator online, will go low when the building AC is transferred to the generator (was Moisture alarm)
- Ch5 Dyersburg Elec./Building AC fail, will go low when Dyersburg Elec. AC lines into transfer switch are dead. (was Intrusion alarm)
- Ch6 Under temp. <50 F (no change)
- Ch7 Over temp. > 90 F (no change)
Steve will cancel phone line.
G and D and the local EMA installed a 25kW generator to power that side of the building. If there's room power-wise would they let us power the rack from it too? We could share the maintenance of the generator. We'd also have to install a transfer switch.
Twisted pair from Richardson is extremely noisy causing problems with
the analog backbone. Richardson is also on the UM 5-year capital
improvement plan to be replaced. We got a tour of Wilder and its doable.
We need to see if we can get a path to CERI so we know whether we just
need to move what's on Richardson now, or whether we need to turn Wilder
into an entirely new node (i.e. new mudworm). Then we can get Planning
to help us with a design and then we can get permission.
We could just move the firewall from 3904/218 to Wilder and get the DBB on the campus internet that way (if there's no aironet line of site). There are good telephone pairs from from the Wilder roof to the basement and we could move the ceri node/mudworm to the basement. Otherwise we'd have to find and trust connections between the basement and CERI.
We have line of site from the southeast corner of 3904 except there are powerlines in the way. Can we put up a tower on the east side of 3904 to get over the trees and power lines on central?
Ownership, Operation, and Maintenance of the four Charleston, SC stations (C1SC, C2SC, ADSC, and TRSC) was turned over the Univ. of SC at Columbia. At Phillip's request, we letting the data flow through CERI for the time being. Eventually it will get switched to Columbia and the network codes will then change from NM to CO.
RDST, CBHT, and MCAR are coming in via gcf2ew on anssworm. CBHT occassionally loses its connection but is overall, ok. MCAR loses its connection about twice as much so while its mostly functional, it should be better. At RDST, a datapacket sneaks through a couple times per day.
Replace foam rubber plugs at all year 1 stations (need to notch it to accomodate the GPS pods): ASAR, CBHT, CUET, CVTN, HDBT, LNXT, MCAR, NAIT, SFTN & TUMT.
NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.
CVVA. John contacted DMME again. He's also looking into the possibility of a freewave link to UVA.
We need a single document (access database excel spreadsheet) to document information on all freefield strongmotion stations to include contact(s), current IP's(eternal/external), what ports are allowed in from what IP's here at CERI, telnet yes or no, trigger levels etc. etc.
Bad CMG5 E/W.
Not getting broadband feed.
Replace two bad mss100's.
We thought the GTTN move was delayed but Chris got an email saying the new tower is done. If so, we need to move our stuff sooner rather than later. Requires at least a two-person crew.
Dynamic ETDown list
CMGA -- roughly daily dropouts (signal fades then goes flat) probably not solar related. Note the counts in the image, it looks like its extremely high gain and going off the rails (on a crazy train?).
GMG - no GPS pod
Gain Channel Problems
ETT -- drifting
GRBT -- spikes on vertical
RBNC.EHE - see image
Strong Motion stations down
CVVA - no internet
GMG- no GPS pod
Only the DYTN upgrade remains.
Agreement is complete: clear to launch.
BHT move. John has a verbal permission from the Catoosa WMA to move it about 9 miles to the E NE to Pilot Knob Mtn. This likely will need to be a written permission. The new location does not have a path to Star Mountain but we do now have hardware for the digital upgrade (a Basalt).
Central Receive: Still have interference from the ham operator. John will pursue this with the FCC.
TRYN: New tower and building are in place, O.K. to install anytime.
WSNC: Fixed transmitter but had to pull the sensor to repair the E-W channel.
CMGA: need to camp out with a spectrum analyzer to identify the interference.
MTTN Node ASTN: Change to AC power. We have permission and can install anytime.
PTGV & PTGX have been down for 6 days ans of 05/10/12.
RLAP, CVMS, HCES, PIGT are out as of 04/05/12. Fixed all but Pigt!! PIGT internet provider has changed his system and knows we need service. Waiting to hear from him,, will keep sending him messages.
RLAP is on a temp. solar panel! And was working when I left. 3 solar regs have failed there???
STLE landowner has asked that we move eq. cabinet. Chris is looking for options. They have not done any building yet (04/04/12).
LCHS and HCES will get upgrades to NetR9. *** Done, having problems with HCES, plan to swap back to the NetRs that was there next week. Also plan to run the conduit. HCES now has the NetRs again and HCEX is back working.
We need a better plan for end-to-end O&M of the GPS network in the post JPD era.
The VLF power supplies were installed CJTR and PIGT. Still needed at NMKM and RLAP.
Chris is trying to find out if we have them or if we need to purchase.
Need to recover the wireless hardware.
Granites are on a 192 network. Steve is setting up a worm to pull data over the wireless and do an end-to-end test for two weeks prior to installation.
The rotation sensors leaked. PMD no longer makes them and isn't willing to repair or replace. New sensors are not in the budget.
HDBT enclosure is reinstalled.
There is about 50' of water in the 100' borehole. The sensor package leaked and is now trashed. It is not repairable and there's no budget for a new sensor.
CMG5TD is ready for installation at HDAR. Need to reinstall hardware for link to AZTN.
Six of the 6TD's will be used for NELE leaving 5 for aftershocks.
Chris is looking for suitable bags to protect them during burial. First experiment with a bag under water failed.
What equipment in the FDL is no longer needed. Mitch will call a meeting of FDL users to discuss culling the equipment. Also, the check out/in process is largely ignored--what system could be implemented that people will actually use. Jer Ming is having trouble with some of the old equipment.
We have one bad CMG-6 Should it go Guralp for repair? Who pays for it?
Done: MORT & LEPT , TNMT & MIST , WADM & STAM , LFRT & GLST , RDGT
Planned upgrades in blue.
All hardware is now on a 4s gain range delay. That is, once it gain ranges, its will stay low gain for at least 4 seconds before going back to high gain. Each cycle that exceeds the threshold resets the counter. So it will return to high gain 4 seconds after the last peak that exceeded the threshold.
Need a list here that shows planned and completed ISIS3.
NSMP Station List.
Newport needs a phoneline. Done and working.
Obion has a gps problem, plan to be in area next week. Done, had a bad antenna.
Paris Tn plan a visit whem looking for Middle tn Site.
Tiptonville Tn hope to visit next week. Done and now on dial up/Phone line.
40T's are obsolesced.
New door locking scheme got delayed to June.
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.